# On convergence structures in graphs

### 09 Mar 2026 v2

@@ Introduction
In the literature, the definition of convergence spaces is more commonly given via filters, but in this work we adopt the net approach, as in [24]. The use of nets in this paper is motivated by convenience: certain arguments can be expressed more directly in terms of eventual behavior along a net. Since filters and nets determine the same convergence structures and can be canonically translated into one another (see [32]), this choice does not affect the underlying theory, but merely streamlines parts of the exposition. Moreover, the formulation in terms of nets often provides more intuition about convergence, as it is typically more natural to think in terms of sequences than in terms of filters.
In this work, we consider the convergence on the vertex set of a graph defined in [9, 31]. Several known results are reproved using nets, and a number of new results concerning the interaction between combinatorial structure of a graph and convergence are established. Whenever a result already exists in the literature, it will be cited in the text; otherwise, it is a new result. As the interaction between graph theory and convergence theory is still relatively unexplored, we expect that this approach will provide a convenient reference point for further research.
@@ 2. Graphs as pretopological spaces
We introduce a natural notion of convergence for nets in a graph, based on the local neighbourhood structure. Let $G$ be a graph and let $\varphi\in\textsc{Nets}(V(G))$ be a net in $V(G)$. We say that $\varphi$ converges to a vertex $v\in V(G)$ if $N[v]\in\varphi^{\uparrow}$. This provides a natural notion of convergence, since the closest one can approach a vertex in a graph is through its neighbours, parametrized by a net. Moreover, as illustrated in Figure 1, a ray may converge to a vertex that is adjacent to every vertex of the ray.
@@ 3. Connectedness and Compactness: Characterizations and Consequences
    Recall that a graph $G$ is said to be 2-connected if it is connected and remains connected after the removal of any single vertex. By the previous theorem, $G$ is $2$-connected if and only if it is connected and, for every $v\in V(G)$, every continuous function $V(G)\setminus\{v\}\to\{0,1\}$ admits a continuous extension to $V(G)$.
@@
To obtain a definition analogous to open covers in the topological case, we need to generalize the concept of open covers. In our setting, these are the convergence systems. Let $\langle X,L\rangle$ be a preconvergence space. A family $\mathcal{C}$ of subsets of $X$ is a local convergence system at $x\in X$ if, for every net $\varphi\in\textsc{Nets}(X)$ such that $\varphi\to_{L}x$, there exists $C\in\mathcal{C}$ with $C\in\varphi^{\uparrow}$. We say that $\mathcal{C}$ is a convergence system if it is a local convergence system for each $x\in X$. The usual terminology for this concept is covering system, as in [2].
@@
Conversely, suppose that $G$ has a finite dominating subset $D$. It is well known that, by Zorn’s Lemma, every filter is contained in an ultrafilter (see [23, Lemma 7.5]). Then it is sufficient to show that every net $\varphi\in\textsc{Nets}(V(G))$ whose induced filter $\varphi^{\uparrow}$ is an ultrafilter converges. Notice that $V(G)=N[D]\in\varphi^{\uparrow}$. Suppose, for the sake of contradiction, that $\varphi$ does not converge. For every $v\in V(G)$ we have $N[v]\notin\varphi^{\uparrow}$. Since $\varphi^{\uparrow}$ is an ultrafilter, it follows that $V(G)\setminus N[v]\in\varphi^{\uparrow}$. In this case, we have $\bigcap_{v\in D}(V(G)\setminus N[v])=V(G)\setminus N[D]=\emptyset\in\varphi^{\uparrow}$, since $D$ is finite. This contradiction tells us that $\varphi$ converges. Therefore, $G$ is compact as convergence space. ∎
@@
    Since the end space and the edge-end space coincide in locally finite graphs, it follows that a compact graph which is locally finite have only finitely many ends. Figure 3 shows that the converse of Corollary 3.3 fails even for locally finite graphs.

### 14 Mar 2026 v3

@@ Introduction
In the literature, the definition of convergence spaces is more commonly given via filters, but in this work we adopt the net approach, as in [26, 35]. The use of nets in this paper is motivated by convenience: certain arguments can be expressed more directly in terms of eventual behavior along a net. Since filters and nets determine the same convergence structures and can be canonically translated into one another (see [34]), this choice does not affect the underlying theory, but merely streamlines parts of the exposition. Moreover, the formulation in terms of nets often provides more intuition about convergence, as it is typically more natural to think in terms of sequences than in terms of filters.
In this work, we consider the convergence on the vertex set of a graph defined in [9, 33]. Several known results are reproved using nets, and a number of new results concerning the interaction between combinatorial structure of a graph and convergence are established. Whenever a result already exists in the literature, it will be cited in the text; otherwise, it is a new result. As the interaction between graph theory and convergence theory is still relatively unexplored, we expect that this approach will provide a convenient reference point for further research. Our approach is similar to that of [18], where the authors relate topological properties of ultrametric spaces to combinatorial properties of labeled trees. This relationship is made possible by the categorical equivalence between ultrametric spaces and trees (see [24]).
@@ 2. Graphs as pretopological spaces
We introduce a natural notion of convergence for nets in a graph, based on the local neighbourhood structure. Let $G$ be a graph and let $\varphi\in\textsc{Nets}(V(G))$ be a net in $V(G)$. We say that $\varphi$ converges to a vertex $v\in V(G)$ if $N[v]\in\varphi^{\uparrow}$. This provides a natural notion of convergence, since the closest one can approach a vertex in a graph is through its neighbours, parametrized by a net. Moreover, as illustrated in Figure 1, a ray (see the definition in Section 3) may converge to a vertex that is adjacent to every vertex of the ray.
@@ 3. Connectedness and Compactness: Characterizations and Consequences
    Recall that a graph $G$ is said to be 2-connected if it is connected and remains connected after the removal of any single vertex. By Theorem 3.1, $G$ is $2$-connected if and only if it is connected and, for every $v\in V(G)$, every continuous function $V(G)\setminus\{v\}\to\{0,1\}$ admits a continuous extension to $V(G)$.
@@
To obtain a definition analogous to open covers in the topological case, we need to generalize the concept of open covers. In our setting, these are the convergence systems. Let $\langle X,L\rangle$ be a preconvergence space. A family $\mathcal{C}$ of subsets of $X$ is a local convergence system at $x\in X$ if, for every net $\varphi\in\textsc{Nets}(X)$ such that $\varphi\to_{L}x$, there exists $C\in\mathcal{C}$ with $C\in\varphi^{\uparrow}$. We say that $\mathcal{C}$ is a convergence system if it is a local convergence system for each $x\in X$. The usual terminology is covering system, as in [2].
@@
Conversely, suppose that $G$ has a finite dominating subset $D$. It is well known that, by Zorn’s Lemma, every filter is contained in an ultrafilter (see [25, Lemma 7.5]). Then it is sufficient to show that every net $\varphi\in\textsc{Nets}(V(G))$ whose induced filter $\varphi^{\uparrow}$ is an ultrafilter converges. Notice that $V(G)=N[D]\in\varphi^{\uparrow}$. Suppose, for the sake of contradiction, that $\varphi$ does not converge. For every $v\in V(G)$ we have $N[v]\notin\varphi^{\uparrow}$. Since $\varphi^{\uparrow}$ is an ultrafilter, it follows that $V(G)\setminus N[v]\in\varphi^{\uparrow}$. In this case, we have $\bigcap_{v\in D}(V(G)\setminus N[v])=V(G)\setminus N[D]=\emptyset\in\varphi^{\uparrow}$, since $D$ is finite. This contradiction tells us that $\varphi$ converges. Therefore, $G$ is compact as convergence space. ∎
@@
    The Figure 3 shows that the converse of Corollary 3.3 fails.